OpenStudy (anonymous):

Solve (x) dy/dx = y + 2x^2y , y(1) = e

OpenStudy (mr.math):

This is a separable differential equation. \[x{dy \over dx}=y(1+2x^2) \implies {dy \over y}={1+2x^2 \over x}dx.\] Now integrate both sides.

OpenStudy (mr.math):

Can you take it from here?

OpenStudy (anonymous):

not really, trying to figure out

OpenStudy (mr.math):

Let me help you with one more step \[\int\limits {dy \over y}=\int\limits ({1 \over x}+2x)dx\] Looks better?

OpenStudy (anonymous):

yeah ! thanks a lot !

OpenStudy (mr.math):

You're welcome.